# Chiara Arina Centre for Cosmology, Particle Physics and Phenomenology (CP3), Université catholique de Louvain, B-1348 Louvain-la-Neuve, Belgium

### v1

@@ III.1.3 Second generation simplified models
Figure 8: NLO/LO ratios of parton luminosities for the NNPDF40 set of parton densities. We consider first the charm-gluon (top left) and strange-gluon (top right) initial states relevant for the $XY$ process, and next the charm-anticharm (bottom left) and strange-antistrange (bottom right) initial states relevant for the $Y\bar{Y}_{t}$ process. Band thickness refers to PDF uncertainties.
@@ III.1.4 Third generation simplified models
These features are illustrated in figure 11, which presents exclusion contours for three classes of third-generation simplified models with fermionic dark matter in the $(M_{Y},M_{X})$ plane. First, the left panel of the figure focuses on supersymmetry-like top-philic S3M_tR scenarios, where the mediator is a coloured scalar similar to a top squark, the dark matter is a Majorana fermion, and the new physics coupling is dynamically set to ensure $\Gamma_{Y}/M_{Y}=0.05$. Since the dependence on the coupling is mild due to the fact that the signal is dominated by its $\lambda$-independent QCD component, using a fixed coupling value of $\lambda=2.2$ as motivated by cosmology would not significantly alter the results. The corresponding figure is therefore omitted from this report. The constraints obtained are relatively weak. The LHC Run 2 is indeed sensitive only to scenarios with mediator masses below $800$ GeV and dark matter masses lighter than about $200$ GeV. These weaker bounds, compared to those for other scenarios with Majorana DM, stem from the distinct decay patterns of the top quarks produced in mediator decays. The latter hence lead to a variety of final states differing from the simpler (hard-scattering-level) one-jet $+E_{T}^{\mathrm{miss}}$ or two-jet $+E_{T}^{\mathrm{miss}}$ signatures. In principle, the limits could be significantly improved by incorporating dedicated $t\bar{t}+E_{T}^{\mathrm{miss}}$ searches in the analysis, as suggested in past studies of composite constructions [27, 28]. However, the absence of validated implementations of such full Run 2 searches in public recasting tools limits this possibility. While we could extrapolate predictions from partial Run 2 results available for some time [171], it is unclear whether such a naive extrapolation would provide meaningful new insights beyond the jet+$E_{T}^{\mathrm{miss}}$ searches already considered, as pointed out in [28]. We therefore refrain from doing so in this report. Finally, as anticipated, the exclusion limits are entirely driven by the $YY_{\mathrm{QCD}}$ channel. While $XX$ contributions could in principle play a role, they are found being negligible and only relevant for scenarios with a mass only slightly larger than that of the top quark.
@@
The remaining panels in figure 12 explore bottom-philic setups. As with the S3M_bR scenarios, we distinguish between cases where the coupling $\lambda$ is fixed based on cosmological considerations (central figures) and those where $\Gamma_{Y}/M_{Y}=0.05$ (rightmost figures). For the fixed coupling case, we use $\lambda=3.8$ for F3S_bR scenarios and $\lambda=0.9$ for F3V_bR scenarios. Again, no new features emerge compared to scalar mediator cases. The limits remain dominated by the $YY_{\mathrm{QCD}}$ channel, with sub-leading contributions from $XY$ production for spectrum featuring light dark matter, and from $YY_{t}$ production for more compressed spectra. Subsequently, mediator masses are excluded up to 1.51.51.51.51.61.61.61.6 TeV in F3S_bR scenarios with dark matter masses below approximately 600600600600700700700700 GeV. Strong constraints are also obtained for compressed spectra in cases where $\lambda$ is derived from $\Gamma_{Y}/M_{Y}=0.05$, but these correspond to baroque setups where perturbative methods are unreliable, so that predictions could not be trusted. For the F3V_bR model, the bounds are similar but slightly weaker, reflecting the use of LO predictions without (differential) $K$-factors.
@@ III.1.5 Simplified models: considerations on signal modelling
Figure 13: Differential distributions for the missing transverse energy $E_{T}^{\mathrm{miss}}$ in the case of the S3M_uR (top row) and F3S_uR (bottom row) scenarios, with mass configurations $\{M_{Y},M_{X}\}=\{3000,10\}$ GeV (left column) and $\{M_{Y},M_{X}\}=\{3000,10\}$ GeV (right column). The distributions represent the number of signal events expected for an integrated luminosity of 300 fb-1 of proton-proton collisions at 13 TeV, and have been obtained using the SFS detector simulation module included in MadAnalysis 5 with the default ATLAS configuration. The contributions of each channel with a cross section larger than 1 ab at both NLO and LO are shown, together with the resulting combined distribution.
In figure 13, we consider the S3M_uR and F3S_uR scenarios, and we show the relative contributions of the various production channels to the $E_{T}^{\mathrm{miss}}$ distribution for two different mass configurations close to the recast bounds at NLO: $M_{Y}=3$ TeV and $M_{X}=10$ or $1500$ GeV. In both cases, the total width of the mediator is set to 5% of its mass (see figure 1), with the $\lambda$ coupling determined accordingly. The differing mass gaps between the mediator and the DM result in significantly distinct shapes for the combined distributions. As already pointed out in [40], at both LO and NLO the $XX$ contribution dominates in the low $E_{T}^{\mathrm{miss}}$ region, but the size of its contribution decreases rapidly with increasing $E_{T}^{\mathrm{miss}}$ values, making other contributions relatively more relevant. In the S3M_uR setup with a large mass gap, the $XY$ contribution becomes dominant around $E_{T}^{\mathrm{miss}}\sim 1$ TeV, peaks between $1.3-1.4$ TeV, and then decreases with less than one event per 25 GeV bin above approximately $1.5$ TeV for an integrated luminosity of $300\,\text{fb}^{-1}$. In contrast, for the F3S_uR scenario, the PDF-enhanced $YY_{t}$ contribution (in which same-charge mediators are produced via $t$-channel DM exchange) dominates above 600 GeV and strongly contributes to shape the $E_{T}^{\mathrm{miss}}$ distribution in the high $E_{T}^{\mathrm{miss}}$ range, as described in [25] (see also, e.g.,  [15]). When the DM state is half the mediator mass, the $YY_{t}$ contribution for S3M_uR again dominates, but this time already for $E_{T}^{\mathrm{miss}}\sim 200$ GeV, with a peak around $1$ TeV, and a fall below one event per bin above $1.7$ TeV for an integrated luminosity of $300\,\text{fb}^{-1}$. In contrast, for the F3S_uR setup, $YY_{t}$ is always, and by far, the dominant contribution.
@@ III.2.1 Flavoured dark matter at the LHC
The case of Dirac flavoured DM coupling to right-handed down-type quarks has been explored in [26]. The strongest LHC constraints arise from searches for supersymmetric squarks in final states with either two light jets or two $b$-tagged jets, both accompanied by missing transverse energy. Additionally, the compressed region is constrained by monojet searches. On the other hand, [74, 73, 176] examined scenarios with couplings to right-handed up-type quarks. Here, the most stringent bounds stem from recasts of supersymmetric squark searches in the $jj+E_{T}^{\mathrm{miss}}$ and $t\bar{t}+E_{T}^{\mathrm{miss}}$ final states, with the precise exclusion limits depending on the flavour structure of the matrix $\lambda$. The blue shaded regions in figure 17 show the resulting constraints, derived from the experimental search [136], on the mediator mass $M_{Y}$ as a function of the first-generation coupling strength $D_{1}$. The limits are shown for two benchmark DM freeze-out scenarios, a first one featuring quasi-degenerate freeze-out (QDF) and a second one predicting single-flavour freeze-out (SFF), their cosmology being discussed in section IV.4.2.1. In addition to these flavour-conserving final states, flavoured DM models also give rise to flavour-violating signatures. For instance, [176] proposed search strategies dedicated to single-top final states, which arise in flavoured DM models coupling to up-type quarks and where mediator pair-production can lead to final states such as $tj+E_{T}^{\mathrm{miss}}$. Similarly, the monotop signature [177], $t+E_{T}^{\mathrm{miss}}$, can also arise, as illustrated in the Feynman diagrams in figure 18. The green and magenta lines in figure 17 show that these search strategies significantly extend the LHC reach beyond the limits obtained from the flavour-conserving channels. Notably, in the case of the SFF benchmark, the monotop search could probe the thermal freeze-out hypothesis, as indicated by the orange dash-dotted line. Finally, Dirac flavoured DM coupling to left-handed quarks has been investigated in [75, 176]. Here, constraints from $jj+E_{T}^{\mathrm{miss}}$ are significantly stronger due to the larger multiplicity of possible partonic final states. Nevertheless, the LHC reach can be further extended through dedicated searches for single-top final states [176]. Interestingly, this model predicts a unique $tb+E_{T}^{\mathrm{miss}}$ signature for which the dedicated search strategy developed in [176] was found to only have limited sensitivity.
Figure 19: Predictions for the single-top charge asymmetry $a_{tj}$ in a model of Majorana flavoured DM coupling to the three flavours of right-handed up-type quarks. We present predictions for $a_{tj}$ (colour-coded) at the 14 TeV LHC, compared with current constraints from tops$+E_{T}^{\mathrm{miss}}$ (light grey) and jets$+E_{T}^{\mathrm{miss}}$ (dark grey) searches, for $D_{2}=0$, $M_{Y}=1200$ GeV, and $M_{X}=400$ GeV (left); we also show the correlations between predictions for $a_{tj}$ and the total single-top cross section $\sigma_{\text{tot}}$ as defined in the denominator of (29), for viable scenarios featuring canonical DM freeze-out (green) and conversion-driven freeze-out (blue). Here, $D_{i}$ represent the DM coupling strengths to the $i^{\mathrm{th}}$ generation, and the figures have been adapted from [77].
@@ III.2.3 Frustrated dark matter at the LHC
In this section, we continue with our study of non-minimal models of $t$-channel dark matter. In the fDM model introduced in section II.4, the only efficient mechanism for DM annihilation involves the $t$-channel exchange of the fermionic mediator $\psi$, where DM particles annihilate into a pair of scalar mediators, $X\bar{X}\to\varphi^{\dagger}\varphi\to qq\bar{q}\bar{q}$. However, this $2\to 4$ process, while relevant for the DM relic abundance, is not directly relevant for the LHC. Instead, the model can be probed at colliders in a number of ways, ranging from jets $+E_{T}^{\mathrm{miss}}$ analyses to direct searches for the colour-charged mediators. The latter category is particularly versatile because colour-charged particles appear in a wide variety of constructions beyond the SM, so many existing searches could be used to constrain this fDM model through reinterpretations. One promising channel is the $t$-channel single production of the scalar mediator $\varphi$ in association with a $W$ boson, as illustrated by the representative Feynman diagram of figure 21 (that includes specific $W$ and $\varphi$ decays). Whereas this $t$-channel process leads to diverse final states depending on the decays of the $\varphi$ mediator and the $W$ boson, the figure illustrates a scenario where the $W$ boson decays hadronically, and the scalar sextet mediator decays into like-sign top quarks ($\varphi\to tt$).
@@
Initial $1.00\times 10^{5}$ $1.00\times 10^{5}$ $1.00\times 10^{5}$
@@ III.2.4 A fermionic portal to a non-Abelian dark sector
For the present analysis, we consider a scenario where the new fermions are partners of the top quark. LHC bounds have been derived by confronting the signal originating from $t_{D}$ pair production, followed by decay into the DM state $V_{D}$ and top quarks, against the results of CMS searches for top squark pair production decaying into DM and tops with partial Run 2 data [196]. Moreover, we have imposed limits on $T\bar{T}$ production, which are approximately given by $m_{T}>1.5$ TeV [197, 198], keeping in mind that single $T$ production is less constrained due to rate suppression by the small $T-t$ mixing. Additionally, we have estimated the relevance of $V^{\prime}$ pair production as well as that of the associated production of a $V^{\prime}$ boson with the Higgs boson, which occur at LO via fermion loops. Representative Feynman diagrams for all these processes are displayed in the left panel of figure 25.
@@ III.3.1 Generalities
In contrast, the superWIMP and freeze-in production mechanisms, which belong to the class of non-thermalised DM scenarios, do not require small mass splittings. Instead, they exhibit extremely large mass hierarchies between the $Y$ and $X$ states in the bulk of the cosmologically viable parameter space [202, 203, 204]. The DM particle in these scenarios can be as light as $\mathcal{O}(10)\,\text{keV}$ [205, 206, 204], effectively rendering it effectively massless with respect to a typical collider scale in large regions of the parameter space. On the other hand, testable mediator masses at the LHC are typically around the TeV scale. Nevertheless, the full cosmologically viable range of mediator masses extends far beyond the LHC reach, leaving only a small fraction of the parameter space accessible to collider searches.
@@ III.3.2 Current LLP searches
In addition, ATLAS has recently performed a search for mildly displaced soft tracks [213]. It relies on a trigger on the missing transverse energy and focuses on events with soft tracks (2 $2\mbox{ GeV}<p_{T}<5\mbox{ GeV}$) exhibiting small displacements ($|d_{0}|<10$ mm). This search specifically targets supersymmetric scenarios with compressed mass spectra, such as setups in which charginos decay to neutralinos and soft pions with small displacements. Despite being a challenging signal due to the presence of the soft objects, the search exploits the relatively small associated SM background rates (i.e. related to the production of soft and displaced pions). It thus allows to look for an LLP signal and hence cover previously existing gaps in typical LLP parameter spaces where $\Delta m<1$ GeV and $c\tau\sim 0.1\mbox{ mm}-1\mbox{ cm}$.
@@ III.3.3.1 Freeze-in/superWIMP regime
[…]
@@ III.3.3.2 Conversion-driven freeze-out regime
Within a quark-philic framework where the DM interacts with right-handed bottom quarks ($f=b$), different LLP searches are sensitive to specific ranges of mediator decay lengths. These include searches for heavy stable charged particles, disappearing tracks and displaced vertices, as well as searches for signals with missing transverse energy whose sensitivity could extend to longer lifetimes depending on their inclusiveness. Let us note that scenarios where $f=u,d,s,c$ yield a very similar phenomenology, but with a different location of the two-body threshold $\Delta m=M_{f}$. In contrast, for $f=t$, the mediator is effectively detector-stable throughout the entire CDFO parameter space as $\Delta m<M_{t}$, making HSCP searches alone well-suited to probing this case [201]. Each class of searches covers different regions of the CDFO parameter space, as shown in figure 32 which illustrates the constraints that could be imposed on the mediator mass and decay length for scenarios compliant with $\Omega h^{2}=0.12$. Bounds from CMS disappearing track searches [223, 243] (green), the ATLAS displaced vertex$+E_{T}^{\mathrm{miss}}$ search [212] (red), and the CMS multijet+$E_{T}^{\mathrm{miss}}$ search [137] (blue) are indicated in the figure, while those from HSCP searches are not shown. While these searches dominate at large $c\tau\gtrsim 10$ m or $\Delta m\lesssim M_{b}$, the corresponding region in parameter space is indeed omitted from the figure. In the detector-stable limit, scenarios with mediator masses up to approximately $1.3$ TeV would however turn out to be constrained [241, 207, 208].
@@
[…]

### v2

@@ III.1.3 Second generation simplified models
Figure 8: NLO/LO ratios of parton luminosities for the NNPDF40 set of parton densities. We consider first the charm-gluon (top left) and strange-gluon (top right) initial states relevant for the XYXY process, and next the charm-anticharm (bottom left) and strange-antistrange (bottom right) initial states relevant for the $Y\bar{Y}_{t}$ process. Band thickness refers to PDF uncertainties.
@@ III.1.4 Third generation simplified models
[…]
@@
The remaining panels in figure 12 explore bottom-philic setups. As with the S3M_bR scenarios, we distinguish between cases where the coupling $\lambda$ is fixed based on cosmological considerations (central figures) and those where $\Gamma_{Y}/M_{Y}=0.05$ (rightmost figures). For the fixed coupling case, we use $\lambda=3.8$ for F3S_bR scenarios and $\lambda=0.9$ for F3V_bR scenarios. Again, no new features emerge compared to scalar mediator cases. The limits remain dominated by the $YY_{\mathrm{QCD}}$ channel, with sub-leading contributions from $XY$ production for spectrum featuring light dark matter, and from $YY_{t}$ production for more compressed spectra. Subsequently, mediator masses are excluded up to 1.51.51.61.6 TeV in F3S_bR scenarios with dark matter masses below approximately 600600700700 GeV. Strong constraints are also obtained for compressed spectra in cases where $\lambda$ is derived from $\Gamma_{Y}/M_{Y}=0.05$, but these correspond to baroque setups where perturbative methods are unreliable, so that predictions could not be trusted. For the F3V_bR model, the bounds are similar but slightly weaker, reflecting the use of LO predictions without (differential) $K$-factors.
@@ III.1.5 Simplified models: considerations on signal modelling
Figure 13: Differential distributions for the missing transverse energy $E_{T}^{\mathrm{miss}}$ in the case of the S3M_uR (top row) and F3S_uR (bottom row) scenarios, with mass configurations $\{M_{Y},M_{X}\}=\{3000,10\}$ GeV (left column) and $\{M_{Y},M_{X}\}=\{3000,1500\}$ GeV (right column). The distributions represent the number of signal events expected for an integrated luminosity of 300 fb-1 of proton-proton collisions at 13 TeV, and have been obtained using the SFS detector simulation module included in MadAnalysis 5 with the default ATLAS configuration. The contributions of each channel with a cross section larger than 1 ab at both NLO and LO are shown, together with the resulting combined distribution.
In figure 13, we consider the S3M_uR and F3S_uR scenarios, and we show the relative contributions of the various production channels to the $E_{T}^{\mathrm{miss}}$ distribution for two different mass configurations close to the recast bounds at NLO: $M_{Y}=3$ TeV and $M_{X}=10$ or $1500$ GeV. In both cases, the total width of the mediator is set to 5% of its mass (see figure 1), with the $\lambda$ coupling determined accordingly. The differing mass gaps between the mediator and the DM result in significantly distinct shapes for the combined distributions. As already pointed out in [40], at both LO and NLO the $XX$ contribution dominates in the low $E_{T}^{\mathrm{miss}}$ region, but the size of its contribution decreases rapidly with increasing $E_{T}^{\mathrm{miss}}$ values, making other contributions relatively more relevant. In the S3M_uR setup with a large mass gap, the $XY$ contribution becomes dominant around $E_{T}^{\mathrm{miss}}\sim 1$ TeV, peaks between 1.31.41.3-1.4 TeV, and then decreases with less than one event per 25 GeV bin above approximately $1.5$ TeV for an integrated luminosity of $300\,\text{fb}^{-1}$. In contrast, for the F3S_uR scenario, the PDF-enhanced $YY_{t}$ contribution (in which same-charge mediators are produced via $t$-channel DM exchange) dominates above 600 GeV and strongly contributes to shape the $E_{T}^{\mathrm{miss}}$ distribution in the high $E_{T}^{\mathrm{miss}}$ range, as described in [25] (see also, e.g.,  [15]). When the DM state is half the mediator mass, the $YY_{t}$ contribution for S3M_uR again dominates, but this time already for $E_{T}^{\mathrm{miss}}\sim 200$ GeV, with a peak around $1$ TeV, and a fall below one event per bin above $1.7$ TeV for an integrated luminosity of $300\,\text{fb}^{-1}$. In contrast, for the F3S_uR setup, $YY_{t}$ is always, and by far, the dominant contribution.
@@ III.2.1 Flavoured dark matter at the LHC
The case of Dirac flavoured DM coupling to right-handed down-type quarks has been explored in [26]. The strongest LHC constraints arise from searches for supersymmetric squarks in final states with either two light jets or two $b$-tagged jets, both accompanied by missing transverse energy. Additionally, the compressed region is constrained by monojet searches. On the other hand, [74, 73, 176] examined scenarios with couplings to right-handed up-type quarks. Here, the most stringent bounds stem from recasts of supersymmetric squark searches in the $jj+E_{T}^{\mathrm{miss}}$ and $t\bar{t}+E_{T}^{\mathrm{miss}}$ final states, with the precise exclusion limits depending on the flavour structure of the matrix $\lambda$. The blue shaded regions in figure 17 show the resulting constraints, derived from the experimental search [136], on the mediator mass $M_{Y}$ as a function of the first-generation coupling strength $D_{1}$. The limits are shown for two benchmark DM freeze-out scenarios, a first one featuring quasi-degenerate freeze-out (QDF) and a second one predicting single-flavour freeze-out (SFF), their cosmology being discussed in section IV.4.2.1. QDF setups consist in scenarios in which all three dark flavours are quasi-degenerate, hence participating all together in the thermal freeze-out process through their combined (co-)annihilations. On the other hand, SFF configurations refer to scenarios where only one particle flavour remains thermally active and undergoes freeze-out, while other flavour states are too heavy to contribute. In addition to these flavour-conserving final states, flavoured DM models also give rise to flavour-violating signatures. For instance, [176] proposed search strategies dedicated to single-top final states, which arise in flavoured DM models coupling to up-type quarks and where mediator pair-production can lead to final states such as $tj+E_{T}^{\mathrm{miss}}$. Similarly, the monotop signature [177], $t+E_{T}^{\mathrm{miss}}$, can also arise, as illustrated in the Feynman diagrams in figure 18. The green and magenta lines in figure 17 show that these search strategies significantly extend the LHC reach beyond the limits obtained from the flavour-conserving channels. Notably, in the case of the SFF benchmark, the monotop search could probe the thermal freeze-out hypothesis, as indicated by the orange dash-dotted line. Finally, Dirac flavoured DM coupling to left-handed quarks has been investigated in [75, 176]. Here, constraints from $jj+E_{T}^{\mathrm{miss}}$ are significantly stronger due to the larger multiplicity of possible partonic final states. Nevertheless, the LHC reach can be further extended through dedicated searches for single-top final states [176]. Interestingly, this model predicts a unique $tb+E_{T}^{\mathrm{miss}}$ signature for which the dedicated search strategy developed in [176] was found to only have limited sensitivity.
[…]
@@ III.2.3 Frustrated dark matter at the LHC
In this section, we continue with our study of non-minimal models of $t$-channel dark matter. In the fDM model introduced in section II.4, the only efficient mechanism for DM annihilation involves the $t$-channel exchange of the fermionic mediator $\psi$, where DM particles annihilate into a pair of scalar mediators, $X\bar{X}\to\varphi^{\dagger}\varphi\to qq\bar{q}\bar{q}$. However, this $2\to 4$ process, while relevant for the DM relic abundance, is not directly relevant for the LHC. Instead, the model can be probed at colliders in a number of ways, ranging from jets $E_{T}^{\mathrm{miss}}$ analyses to direct searches for the colour-charged mediators. The latter category is particularly versatile because colour-charged particles appear in a wide variety of constructions beyond the SM, so many existing searches could be used to constrain this fDM model through reinterpretations. One promising channel is the $t$-channel single production of the scalar mediator $\varphi$ in association with a $W$ boson, as illustrated by the representative Feynman diagram of figure 21 (that includes specific $W$ and $\varphi$ decays). Whereas this $t$-channel process leads to diverse final states depending on the decays of the $\varphi$ mediator and the $W$ boson, the figure illustrates a scenario where the $W$ boson decays hadronically, and the scalar sextet mediator decays into like-sign top quarks ($\varphi\to tt$).
@@
Initial $0\times 10^{5} 1.00×1$$0\times 10^{5}$ $1.00\times 10^{5}$
@@ III.2.4 A fermionic portal to a non-Abelian dark sector
For the present analysis, we consider a scenario where the new fermions are partners of the top quark. LHC bounds have been derived by confronting the signal originating from $t_{D}$ pair production, followed by decay into the DM state $V_{D}$ and top quarks, against the results of CMS searches for top squark pair production decaying into DM and tops with partial Run 2 data [196]. Moreover, we have imposed limits on $T\bar{T}$ production, which are approximately given by $m_{T}>1.5$ TeV [197, 198], keeping in mind that single $T$ production is less constrained due to rate suppression by the small TtT-t mixing. Additionally, we have estimated the relevance of $V^{\prime}$ pair production as well as that of the associated production of a $V^{\prime}$ boson with the Higgs boson, which occur at LO via fermion loops. Representative Feynman diagrams for all these processes are displayed in the left panel of figure 25.
@@ III.3.1 Generalities
In contrast, the superWIMP and freeze-in production mechanisms, which belong to the class of non-thermalised DM scenarios, do not require small mass splittings. Instead, they often exhibit extremely large mass hierarchies between the $Y$ and $X$ states in the bulk of the cosmologically viable parameter space [202, 203, 204]. The DM particle in these scenarios can be as light as $\mathcal{O}(10)\,\text{keV}$ [205, 206, 204], effectively rendering it effectively massless with respect to a typical collider scale in large regions of the parameter space. On the other hand, testable mediator masses at the LHC are typically around the TeV scale. Nevertheless, the full cosmologically viable range of mediator masses extends far beyond the LHC reach, leaving only a small fraction of the parameter space accessible to collider searches.
@@ III.3.2 Current LLP searches
In addition, ATLAS has recently performed a search for mildly displaced soft tracks [213]. It relies on a trigger on the missing transverse energy and focuses on events with soft tracks ($2\mbox{ GeV}<p_{T}<5\mbox{ GeV}$) exhibiting small displacements ($|d_{0}|<10$ mm). This search specifically targets supersymmetric scenarios with compressed mass spectra, such as setups in which charginos decay to neutralinos and soft pions with small displacements. Despite being a challenging signal due to the presence of the soft objects, the search exploits the relatively small associated SM background rates (i.e. related to the production of soft and displaced pions). It thus allows to look for an LLP signal and hence cover previously existing gaps in typical LLP parameter spaces where $\Delta m<1$ GeV and $c\tau\sim 0.1\mbox{ mm}-1\mbox{ cm}$.
@@ III.3.3.1 Freeze-in/superWIMP regime
Figure 30: Isolines of constant reheating temperature $T_{\text{rh}}$ expressed in the $(M_{Y},c\tau)$ plane, for scenarios accounting for the observed DM relic abundance and a DM mass of 12 keV for a muon-philic Majorana DM model (left), and 10 keV for a top-philic model with scalar DM (right). In the left panel, adapted from [254], we study the impact of the reheating potential $V(\Phi)\sim\Phi^{k}$, and consider $T_{\text{rh}}=20\,\text{GeV}$ (blue), $100\,\text{GeV}$ (black) and $10^{4}\,\text{GeV}$ (red, with details of reheating being no longer relevant for the considered mediator masses). Solid lines apply to $k=2$, while dashed and dot-dashed lines illustrate fermionic and bosonic reheating scenarios for $k=4$ potentials respectively. Various constraints from LLP searches are additionally shown as grey-shaded regions. In the right panel, adapted from [252], we consider $k=2$, vary $T_{\text{rh}}$, and show collider constraints from DJ (blue, [237]), DV (green, [212, 244]) and DL (dark green, [214]) searches.
@@ III.3.3.2 Conversion-driven freeze-out regime
[…]
@@
We continue our discussion with leptophilic scenarios, that we illustrate with a setup involving a Majorana DM candidate coupled to a scalar slepton-like mediator [260]. For an interaction strength $\lambda\lesssim 10^{-6}$, neglecting bound state formation effects [258, 261] and ignoring Higgs-mediator couplings, the measured DM relic abundance can be achieved through CDFO production for mediator masses below $200~\text{GeV}$ and mass splittings $\Delta m<2.6~\text{GeV}$. While such a setup is already stringently constrained by searches at the LHC, introducing a non-zero quartic coupling between the scalar mediator and the Higgs boson significantly enlarges the viable regions in the model parameter space. In figure 33, we consider such a quartic coupling $\lambda_{H}=0.1$, and illustrate the constraints on the parameter space of a model including interactions with the right-handed muon. The results are as usual given in terms of the mediator lifetime and mass. The upper region of the plot corresponds to long-lived mediators with decay lengths of a few centimetres or more, or equivalently small values of the DM-mediator coupling together with a very compressed mass spectrum (as indicated by the blue contours). For $\lambda_{H}=0.1$, the CDFO regime extends to mediator masses up to $300~\text{GeV}$, and necessitates mass splittings $\Delta m<3~\text{GeV}$. We include in the figure the complementary constraints emerging from early LHC Run 2 searches for HSCP [240, 219] (red) and disappearing tracks [242, 243] (green and purple), that should get a little bit stronger by relying on more recent analyses like [223, 241].